OpenStudy (anonymous):

What is the equation of the line that passes through the point (4,1) and has a slope of 5?

OpenStudy (amistre64):

which form of the equation?

OpenStudy (bahrom7893):

id like y = mx+b amistre

OpenStudy (amistre64):

y = mx =mPx + Py y = 5x -5(4) + 1 is one form

OpenStudy (amistre64):

..... typos lol y = mx -mPx + Py

OpenStudy (anonymous):

a. y=5x-3 b.y=5x-1 c.y=5x+3 d.y=5x-19

OpenStudy (amistre64):

then that should do fine